OpenStudy (anonymous):

a = 20, b = 28, c = ?

OpenStudy (anonymous):

a^2 + b^2 = c^2 20^2 + 28^2 = c^2 400 + 784 = c^2 1184 = c^2 sq rt 1184 = sq rt c^2 34.409 (rounded = 34.41) = c